# Linear fractional relative risk aversion

### v1

@@ Abstract
We characterize the family of utility functions satisfying linear fractional relative risk aversion (LFRRA) in terms of the Gauss hypergeometric functions. We apply this family, which nests various utility functions used in different strands of literature, to monopolistic competition and obtain a closed-form solution for the profit-maximizing price by generalizing the Lambert $W$ function. We let firm-level data decide whether the RRA in each sector or in the aggregate economy is increasing, decreasing, or constant, which in turn determines whether markups are decreasing, increasing, or constant with respect to marginal costs.
@@ Introduction
In these fields of economics, different parametric specifications of $u$ have been used for deriving qualitative and quantitative results. Examples include the utility functions with hyperbolic relative risk aversion (HARA)—which nests both constant relative risk aversion (CRRA) and constant absolute risk aversion (CARA) as special cases—and the constant revenue elasticity of marginal revenue (CREMR) utility function, which also nests CRRA but does not belong to HARA. [^2] See, e.g., Merton (1971) for HARA and Mrázová et al. (2021) for CREMR.
We propose a parametric family of utility functions that not only unifies these existing specifications into a single framework but also allows for new ones. To this end, we start with linear fractional relative risk aversion (LFRRA), defined as $\frac{qu^{\prime\prime}(q)}{u^{\prime}(q)}=\frac{\alpha q+\beta}{\alpha\sigma q+1}$, where $\alpha$, $\beta$, and $\sigma$ are parameters. We then show that this second-order differential equation can be rewritten as a hypergeometric differential equation, $z(1-z)v^{\prime\prime}(z)+(\beta-\frac{z}{\sigma})v^{\prime}(z)=0$, using the change of variables $z=-\alpha\sigma q$ with $u(q)=v(z)$. Solving this differential equation allows us to characterize the family of utility functions satisfying LFRRA in terms of the Gauss hypergeometric functions.
We apply the LFRRA family of utility functions thus obtained to a model of monopolistic competition with heterogeneous firms, where the profit-maximizing price $p$, given marginal cost $m$, is defined as the solution to $\frac{p-m}{p}=-\frac{qu^{\prime\prime}(q)}{u^{\prime}(q)}$. We derive the closed form solution for $p=\frac{m}{\mu}$, where $\mu$ is the inverse markup function parametrized by $\{\alpha,\beta,\sigma\}$. It turns out that $\mu$ generalizes the Lambert $W$ function (Lambert, 1758; Euler, 1783; Corless et al., 1996), which has been used for the inverse markup for the CARA utility function in Behrens et al. (2014, 2017, 2020). Our generalized Lambert $W$ function reduces to the original when $\alpha>0$ and $\beta=\sigma=0$.
@@
The remainder of the paper is organized as follows. Section 2 introduces the LFRRA, characterizes the family of utility functions satisfying the LFRRA using the Gauss hypergeometric functions, and discusses a wide range of special cases. Section 3 embeds the LFRRA family of utility functions into a monopolistic competition model with heterogeneous firms, derives the inverse markup function in terms of the generalized Lambert $W$ function, and provides properties of markups, prices, and quantities. Section 4 takes the LFRRA family to data, provides estimates of our key parameters $\{\alpha,\beta,\sigma\}$, and detects which sectors display increasing, decreasing, or constant RRA. Section 5 shows that our monopolistic competition analysis can be extended to the case of implicitly additive preferences. Finally, Section 6 concludes.
@@
which is the hypergeometric differential equation $z(1-z)v^{\prime\prime}(z)+[c-(a+b+1)z]v^{\prime}(z)-abv(z)=0$ with $a$, $b$, and $c$ satisfying $c=\beta$, $a+b+1=\frac{1}{\sigma}$, and $ab=0$. Letting $c_{1}$ and $c_{2}$ denote arbitrary coefficients, the solution to this hypergeometric differential equation is given by $v(z)=\textstyle c_{1}\times{}_{2}F_{1}(a,b;c;z)+c_{2}\times z^{1-c}{}_{2}F_{1}(1+a-c,1+b-c;2-c;z),$ where $_{2}F_{1}(a,b;c;z)=\sum_{n=0}^{\infty}\frac{(a)_{n}(b)_{n}}{(c)_{n}}\frac{(z)^{n}}{n!}$, with $\lvert z\rvert<1$, denotes the hypergeometric function and $(\cdot)_{n}$ is the (rising) Pochhammer symbol (see Seaborn, 1991). Solving $c=\beta$, $a+b+1=\frac{1}{\sigma}$, and $ab=0$ for $a$, $b$, and $c$ yields two solutions $\{a,b,c\}=\{\frac{1-\sigma}{\sigma},0,\beta\}$ or $\{0,\frac{1-\sigma}{\sigma},\beta\}$ that are symmetric in $a$ and $b$. Without loss of generality, we take the first one, so that the solution to (3) is given by
@@
where $_{2}F_{1}\left(1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\right)=\sum_{n=0}^{\infty}\frac{(1-\frac{1}{\sigma})_{n}(1)_{n}}{(2-\beta)_{n}}\frac{(-\alpha\sigma q)^{n}}{n!}$, with $\lvert-\alpha\sigma q\rvert<1$, denotes the hypergeometric function; $(\cdot)_{n}$ is the (rising) Pochhammer symbol; and $K>0$ and $C$ are constants.

### v2

@@ Abstract
We characterize the family of utility functions satisfying linear fractional relative risk aversion (LFRRA) in terms of the Gauss hypergeometric functions. We apply this family, which nests various utility functions used in different strands of literature, to monopolistic competition and obtain the profit-maximizing price by generalizing the Lambert $W$ function. We let firm-level data decide whether the RRA in each sector or in the aggregate economy is increasing, decreasing, or constant, which in turn determines whether markups are decreasing, increasing, or constant with respect to marginal costs.
@@ Introduction
In these fields of economics, different parametric specifications of $u$ have been used for deriving qualitative and quantitative results. Examples include the utility functions with hyperbolic absolute risk aversion (HARA)—which nests both constant relative risk aversion (CRRA) and constant absolute risk aversion (CARA) as special cases—and the constant revenue elasticity of marginal revenue (CREMR) utility function, which also nests CRRA but does not belong to HARA. [^2] See, e.g., Merton (1971) for HARA and Mrázová et al. (2021) for CREMR.
We propose a parametric family of utility functions that not only unifies these existing specifications into a single framework but also allows for new ones. To this end, we start with linear fractional relative risk aversion (LFRRA), defined as $\frac{qu^{\prime\prime}(q)}{u^{\prime}(q)}=\frac{\alpha q+\beta}{\alpha\sigma q+1}$, where $\alpha$, $\beta$, and $\sigma$ are parameters. We then show that this second-order differential equation can be rewritten as a hypergeometric differential equation, $z(1-z)v^{\prime\prime}(z)+(\beta-\frac{z}{\sigma})v^{\prime}(z)=0$, using the change of variables $z=-\alpha\sigma q$ with $u(q)=v(z)$. Solving this differential equation allows us to characterize the family of utility functions satisfying LFRRA in terms of the Gauss hypergeometric functions (Gauss, 1812).
We apply the LFRRA family of utility functions thus obtained to a model of monopolistic competition with heterogeneous firms, where the profit-maximizing price $p$, given marginal cost $m$, is defined as the solution to $\frac{p-m}{p}=-\frac{qu^{\prime\prime}(q)}{u^{\prime}(q)}$. We derive the solution, $p=\frac{m}{\mu}$, where $\mu$ is the inverse markup function parametrized by $\{\alpha,\beta,\sigma\}$. It turns out that $\mu$ generalizes the Lambert $W$ function (Lambert, 1758; Euler, 1783; Corless et al., 1996), which has been used for the inverse markup for the CARA utility function in Behrens et al. (2014, 2017, 2020). Our generalized Lambert $W$ function reduces to the original when $\alpha>0$ and $\beta=\sigma=0$.
@@
The remainder of the paper is organized as follows. Section 2 introduces the LFRRA, characterizes the family of utility functions satisfying the LFRRA using the Gauss hypergeometric functions, and discusses a wide range of special cases. Section 3 embeds the LFRRA family of utility functions into a monopolistic competition model with heterogeneous firms, derives the inverse markup function in terms of the generalized Lambert $W$ function, and provides properties of markups, prices, and quantities. In Section 4, we extend our monopolistic competition analysis to the case of implicitly additive preferences and generalize the LFRRA family to some cases of non-linear fractional forms. Section 5 takes the LFRRA family to data, provides estimates of our key parameters $\{\alpha,\beta,\sigma\}$, and detects which sectors display increasing, decreasing, or constant RRA. We further compare these results with those obtained from the other specifications discussed in Section 4. Finally, Section 6 concludes.
@@
which is the hypergeometric differential equation $z(1-z)v^{\prime\prime}(z)+[c-(a+b+1)z]v^{\prime}(z)-abv(z)=0$ with $a$, $b$, and $c$ satisfying $c=\beta$, $a+b+1=\frac{1}{\sigma}$, and $ab=0$. Letting $c_{1}$ and $c_{2}$ denote arbitrary coefficients, the solution to this hypergeometric differential equation is given by $v(z)=\textstyle c_{1}\times{}_{2}F_{1}(a,b;c;z)+c_{2}\times z^{1-c}{}_{2}F_{1}(1+a-c,1+b-c;2-c;z),$ where $_{2}F_{1}(a,b;c;z)=\sum_{n=0}^{\infty}\frac{(a)_{n}(b)_{n}}{(c)_{n}}\frac{z^{n}}{n!}$, with $\lvert z\rvert<1$, denotes the Gauss hypergeometric function and $(\cdot)_{n}$ is the (rising) Pochhammer symbol (see Seaborn, 1991). Solving $c=\beta$, $a+b+1=\frac{1}{\sigma}$, and $ab=0$ for $a$, $b$, and $c$ yields two solutions $\{a,b,c\}=\{\frac{1-\sigma}{\sigma},0,\beta\}$ or $\{0,\frac{1-\sigma}{\sigma},\beta\}$ that are symmetric in $a$ and $b$. Without loss of generality, we take the first one, so that the solution to (3) is given by
@@
where $_{2}F_{1}\left(1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\right)=\sum_{n=0}^{\infty}\frac{(1-\frac{1}{\sigma})_{n}(1)_{n}}{(2-\beta)_{n}}\frac{(-\alpha\sigma q)^{n}}{n!}$, with $\lvert-\alpha\sigma q\rvert<1$, denotes the Gauss hypergeometric function; $(\cdot)_{n}$ is the (rising) Pochhammer symbol; and $K>0$ and $C$ are constants.